I have to word problem that I need to have detailed answers for and asap thank you
1.Flight of a Ball: A ball is hit upward. Its height h in feet after t seconds is given by h = - 16tē + 80t +4. At what time is the ball at 100 feet in the air?
2.Revenue: A company makes tops for the boxes of pickup trucks. The total revenue R in dollars from selling the tops for p dollars each is given by R =p(200-p), where p≤ 200.
(a) Find R when p = 100
(b) Find p when R = $7500
(c) If you have a calculator, use it to solve part (b) numerically with a table of values. Do your answers agree?
Please I need help with these two problems asap

Answer by mananth(16949)

1.Flight of a Ball: A ball is hit upward. Its height h in feet after t seconds is given by h = - 16tē + 80t +4. At what time is the ball at 100 feet in the air?
h = - 16tē + 80t +4
h=100
100= -16^2+80t+4
16t^2-80t-4+100=0
16t^2-80t+96=0
/16
t^2-5t+6=0
Factorize
(t-3)(t-2)=0
t=3 OR 2 seconds

2.Revenue: A company makes tops for the boxes of pickup trucks. The total revenue R in dollars from selling the tops for p dollars each is given by R =p(200-p), where p≤ 200.
(a) Find R when p = 100
R=p(200-p)
R=100(200-100)
R=100*100
R= $ 10,000
(b) Find p when R = $7500
7500 = p(200-p)
7500=200p-p^2
p^2-200p+7500=0
p^2-150p-50p+7500=0
p(p-150)-50(p-150)=0
(p-50)(p-150)=0
p= 50 OR 150
(c) If you have a calculator, use it to solve part (b) numerically with a table of values. Do your answers agree?
m.ananth@hotmail.ca